# The physical basis of teleological evolutions

### v3

@@ 2.2 Teleological character of natural evolutions – a possible physical basis
This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Naturally, at the same time we should adopt quantum cosmology $\left[12,13\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[14-18\right]$, the “fine-tuned universe” version of it. This version of the principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”. Let us assume, for simplicity, that this correlation today is one-to-one. The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
@@
By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers). Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life. The measurement of either observable today would select the eigenvalues of both observables. All measurement outcomes are as if the value of the fundamental constants was selected at the beginning of time.
Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants). We have seen that this unitary evolution can always occur with quantum speedup. Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[19\right]$. It should consequently be teleological in character.
@@
That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence. Things become quite different in the case of the evolutions of the living. The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling. It might go along with ancient beliefs, but of a kind deliberately ignored by science. Once science was accused of heresy for denying a physical basis to beliefs of the kind in question. We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
@@ 3 Positioning the work
We try and position the present work. Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Besides the retrocausal approach to the quantum computational speedup $\left[4-8\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
@@
Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
@@ 4 The time-symmetric quantum causal loops – formalization
We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[9\right]$quantum search algorithm. As anticipated in the Rationale, this immediately generalizes first to any number of drawers (under Long’s $\left[10\right]$ version of Grover algorithm that always provides the solution with certainty), then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. The formalization consists of the following steps:
@@ 4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow333 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operantion that changes nothing for us. .
@@
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement444 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution. Knowledge of the problem-setting selected by the initial measurement would tell her also the solution. .
@@ 5 The physical basis of the notion of teleological evolution
We have seen, rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question. In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces. As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
@@
As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm. This search can therefore occur with quantum speedup and does even if it is random. Consequently, the evolution of a universe toward the value of life turns out to be teleological in character. In the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
Eventually, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within a quantum cosmological view.
@@ 5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
Instead, a decohered process endowed with quantum speedup must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process. Let us call time-symmetric also the kind of stochastic evolution that we have just outlined.
@@
The usual explanation is of course a causal one: at first, feathers appeared at random and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
@@ 5.3 How to model a teleological natural evolution
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
@@ 5.5 Impact on philosophy
Of course, the current theory of teleological evolutions rehabilitates an ancient notion of natural philosophy $\left[39\right]$. We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view555 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
@@
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$, and the present theory is a mutual causality view.
@@
For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that what we see around and inside is essentially inseparable from the whole.
@@ 6 Research prospects
(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
(iii) Explore the possibility that other quantum features suspected to be involved in the fundamental processes of life derive top-down from a quantum character of the universe (as, presently, the teleological character of the evolutions of the living).
(iv) Look for the possible impact of the teleological view on quantum cosmology. According to current assumptions, even the inanimate part of it, anyway an evolution toward life, should be endowed with quantum speedup and thus teleological in character.
@@
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
@@ 7 Conclusions
The work is, of course, unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
@@
Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings. We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement. This is a key feature of quantum algorithms, not visible in their customary quantum description666 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting. . We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
@@
Eventually, we should take the quantum superposition of all the causal loops.
@@
If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces. This creates the room for a quantum speedup. There should be a quantum speedup even in the case that the search for one-to-one correlation is random.
@@
This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years. The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
We should now call into play the “fine-tuned universe” version of the Anthropic Principle. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one
@@
Of course, it is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
Regarding quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling. We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption. Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience the strange character of the microscopic quantum world.

### v4

@@ 2.2 Teleological character of natural evolutions – a possible physical basis
This, of course, immediately raises the problem of explaining the existence of quantum features – indeed the quantum speedup – in evolutions that can be the size of the Earth and have a duration of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels suggests the possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Naturally, at the same time we should adopt quantum cosmology $\left[13,14\right]$, namely the assumption that the evolution of the universe, a perfectly isolated system since by definition there is nothing outside it, is quantum and unitary in character.
An immediate consequence is having to also adopt a quantum cosmological interpretation of the Anthropic Principle $\left[15-19\right]$, the “Fine-tuned Universe” version of it. This version of the Principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental physical constants since the beginning of time and the “value” of life today, in particular “sentient life” or “no life”. Let us assume, for simplicity, that this correlation today is one-to-one. The significance of this principle lies in the fact that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
@@
By the way, since the value of the fundamental constants never changes along the universe’s evolution, in the universe we are in we can do without the initial measurement (when there were no observers). Today, there would be a quantum superposition of tensor products, each a possible value of the fundamental constants multiplying the corresponding value of life. The measurement of either observable today would select the eigenvalues of both maximally entangled observables. Everything is as if the value of the fundamental constants was selected at the beginning of time.
Summing up, we are dealing with a unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue selected by the initial measurement (here the value of the fundamental constants). We have seen that this unitary evolution can always occur with quantum speedup – there is the room for it. Being a quantum search (for one-to-one correlation), it should occur with quantum speedup even if the search (by the universe’s evolution under the fundamental physical laws and constants) is random $\left[12\right]$. It should consequently be teleological in character.
@@
That quantum algorithms are teleological evolutions might even be acceptable from the psychological standpoint: it would be the usual strangeness of the microscopic quantum world and its short lived evolutions soon killed by decoherence. Things become quite different in the case of the evolutions of the living. The fact that what happens around and within us would be driven not only by the past but also by a future even billions of years ahead is more than puzzling. It might also go along with ancient beliefs, but of a kind deliberately ignored by science. Once science was accused of heresy for denying a physical basis to beliefs of the kind in question. We could be facing the reverse problem now: proposing a physical basis to the notion of teleological evolution could well be seen as a scientific heresy.
@@ 3 Positioning the work
We try and position the present work. Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Besides the retrocausal approach to the quantum computational speedup $\left[1-5\right]$, we should mention time-symmetric quantum mechanics $\left[20-24\right]$, in particular its study of the possibility that a future choice changes a past quantum measurement outcome $\left[23\right]$, and ongoing attempts to study quantum processes without the need to preassume a definite causal structure $\left[25-28\right]$.
@@
Other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[13\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the value of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating from before Deutsch’s 1985 discovery, the two principles in question make no reference to the quantum speedup of the universe’s evolution and, a-fortiori, its consequent teleological character.
@@ 4 The time-symmetric quantum causal loops – formalization
We formalize the notion of the time-symmetric quantum causal loops on the four drawer instance of Grover $\left[6\right]$ quantum search algorithm. As anticipated in the Rationale, this immediately generalizes first to any number of drawers, then to any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. The formalization consists of the following steps:
@@ 4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow444 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operantion that changes nothing for us. .
@@
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement555 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution. Knowledge of the problem-setting selected by the initial measurement would tell her also the solution. .
@@ 5 The physical basis of the notion of teleological evolution
We have seen, by the way rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question. In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces. As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
@@
As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm. This search can therefore occur with quantum speedup and does even if it is random. Consequently, the evolution of a universe toward the value of life turns out to be teleological in character. Of course, in the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
Eventually, as we have seen, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within the quantum cosmological view.
@@ 5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
Instead, a decohered process endowed with quantum speedup must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process. Since also in this case the time-symmetric causal loops we are dealing with should be involved, let us call time-symmetric also the just outlined kind of stochastic evolution.
@@
The usual explanation is of course a causal one: at first, feathers appeared through random mutations and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
@@ 5.3 How to model a teleological natural evolution
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious (approximate) way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
@@ 5.5 Impact on philosophy
Of course, the current theory of teleological evolutions would rehabilitate an ancient notion of natural philosophy $\left[39\right]$. We should therefore be allowed to look for the possibility that there are other relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
@@
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, the anthropic cosmological principles of Barrow and Tipler $\left[13\right]$, and the present theory is a mutual causality view.
@@
For what concerns reductionism vs monism, we note that deriving the teleological character of the decohered evolutions of the living systems on Earth from that of the evolution of a quantum universe is an essentially monistic view. It presupposes that what we see around and within us is essentially inseparable from the whole.
@@ 6 Research prospects
(ii) Look for toy models that capture the mechanism for which the evolution of life on Earth (the mostly autonomous evolution of a subsystem), although decohered, inherits the quantum speedup and consequently the teleological character of the evolution of the universe (the system) toward life.
(iii) The mechanism described in (ii) might apply to other life processes for which the presence of a quantum behavior is suspected, for example, to the possible quantum behavior of the brain. Although mostly autonomous from the rest of the universe, this behavior could derive from that of a quantum universe by tracing out everything external to the brain. Note the radical difference between this view and the usual one where people looks for an autonomous quantum behavior in the macroscopic, warm and wet brain. By the way, we could introduce a distinction between two kinds of quantum behaviors: autonomous, like those of microscopic systems sufficiently isolated from all the rest, and derived, namely quantum behaviors of even macroscopic subsystems obtained from that of a quantum universe – the system – by tracing out all the rest.
(iv) Look for the possible impact of the teleological view on quantum cosmology. According to current assumptions, even the inanimate part of the latter, always an evolution toward life, should be endowed with quantum speedup and thus have teleological character.
@@
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated as pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
@@ 7 Conclusions
The work is, of course, highly unconventional. It might be useful to summarize in a consequential way the line of reasoning followed.
@@
Thus, as a first step, we extend the customary quantum description of the process of solving the problem to the upstream process of setting the problem, by adding the initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings. We have thus a unitary evolution between two one-to-one correlated measurement outcomes that does not change the problem-setting selected by the initial measurement. This is a key feature of quantum algorithms, not visible in their customary quantum description777 If the solution is not an invertible function of the problem-setting, to have one-to-one correlation between the two measurement outcomes, it suffices that the problem-solver, at the end, measures both the solution and the problem-setting. . We call this kind of process the ordinary quantum description of the quantum algorithm, ordinary because it is still a causal description that cannot describe causal loops.
@@
Eventually, we should take the quantum superposition of all the causal loops. Of course, the time-symmetrization leaves the ordinary quantum description mathematically unaltered but changes its single causal process into a quantum superposition of causal loops.
@@
If the unitary evolution builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement, like in the case of Grover algorithm, in each time-symmetrization instance the dimension of the Hilbert space in which the search for one-to-one correlation takes place reduces. This creates the room for a quantum speedup. There is a quantum speedup even in the case that the search for one-to-one correlation is random.
@@
This immediately raises the problem of explaining a quantum speedup in evolutions of the living that can be the size of the Earth and have durations of billions of years. The fact it is obviously impossible to scale up the quantum features of the microscopic world to such levels suggests the only possible solution: scaling them down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should adopt quantum cosmology, namely the assumption that the universe evolves in a unitary quantum way.
We should now call into play the “Fine-tuned Universe” version of the Anthropic Principle. It asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value of the fundamental physical constants since the beginning of time and the “value” of life today, in particular, either “existence of sentient life” or “absence of life”. We assume for simplicity that this correlation is one-to-one.
@@
Of course, that of teleological evolution is a puzzling notion that radically changes a long-standing way of viewing the role of causality in the physical world.
Regarding the quantum phenomena in the microscopic world, we think to have provided clear evidence that the notion of the time-symmetric quantum causal loops is a necessary consequence of the notion of quantum computational speedup.
Regarding the teleological character of the macroscopic evolutions of life, the situation becomes even more puzzling. We limit ourselves to noticing that this is a necessary consequence of the quantum cosmological assumption. Seeing the macroscopic evolutions of life as the partial trace over everything else of a quantum evolution of the universe toward life brings into the macroscopic world of everyday experience some of the strange characters of the microscopic quantum world.